A cell phone battery manufacturer claims that one of their batteries for a particular cell phone will outperform a competitorA cell phone battery manufacturer claims that one of their batteries for a particular cell phone will outperform a competitorA quality control manager is comparing shipments of android tablets from two suppliers. Identify these suppliers as SupplierA quality control manager is comparing shipments of android tablets from two suppliers. Identify these suppliers as Supplierplaced A company makes two types of batteries (Type A and Type B) for midsize electric cars. These two batteries are the distA company makes two types of batteries (Type A and Type B) for midsize electric cars. These two batteries are the distance (iwas conducted to investigate the proportion of professors in colleges in a particular state who were involved with grant fundThe budget director for a company is interested in comparing the daily travel expenses for the sales staff and the audit stafThe budget director for a company is interested in comparing the daily travel expenses for the sales staff and the audit stafTwo different methods of teaching an elementary statistics course were . Identify these two methods as Method A and Method B.Step 1 of 2: Construct a 90 percent confidence interval for the difference in the average percentages for the two classes (MeTreated Group Control Group =8.3 S1=0.7 n= 45 2 =3.4 S2= 1.3 nz= 45 Step 1 of 2: Construct a 99 percent confidence interval fTreated Group i=8.3 Control Group i2 3.4 S1 = 0.7 n= 45 S2= 1.3 nz= 45 Step 2 of 2: Interpret the correct answer from STEP 1The Lake Eerie Islands of northern Ohio are renowned for their fine wineries. Suppose a vintner of one of the Lake Eerie IslaThe Lake Eerie Islands of northern Ohio are renowned for their fine wineries. Suppose a vintner of one of the Lake Eerie IslaWhen confidence intervals for the difference between two population means are constructed, if the population variances are un

Please do all questions...
Answer #1

1)

step 1)

sample #1   ----->   manufacture
first sample size,     n1=   55
number of successes, sample 1 =     x1=   41
proportion success of sample 1 , p̂1=   x1/n1=   0.7455
      
sample #2   ----->   competitor
second sample size,     n2 =    55
number of successes, sample 2 =     x2 =    44
proportion success of sample 1 , p̂ 2=   x2/n2 =    0.8000

level of significance, α =   0.05              
Z critical value =   Z α/2 =    1.960   [excel function: =normsinv(α/2)      
                  
Std error , SE =    SQRT(p̂1 * (1 - p̂1)/n1 + p̂2 * (1-p̂2)/n2) =     0.0797          
margin of error , E = Z*SE =    1.960   *   0.0797   =   0.1563
                  
confidence interval is                   
lower limit = (p̂1 - p̂2) - E =    -0.055   -   0.1563   =   -0.2108
upper limit = (p̂1 - p̂2) + E =    -0.055   +   0.1563   =   0.1018
                  
so, confidence interval is (   -0.2108   < p1 - p2 <   0.1018   )  

step 2)

we are 95% confident that true difference in proportions of batteries which lasted beyond 2000 hours for manufacture's brand relative to competitor's lies within confidence interval

please note that only one ques per post can be answered as per HomeworkLib answering guidelines